# Dressed basis sets for the modeling of exchange interactions in double quantum dots

### v1

@@ I Introduction
In this work, we introduce a methodology for the efficient modeling of Coulomb interactions in double quantum dot (DQD) devices. Starting from reference CI calculations, we introduce a reduced basis set of “dressed” states accounting for the main Coulomb correlations. This basis set can describe the lowest singlet and triplet states relevant for exchange-driven operations over the whole operational gate voltages range with as few as a hundred basis functions (whereas the original CI basis set contains $>4500$ Slater determinants). This allows for fast explorations of the device physics and for efficient time-dependent simulations. We illustrate this framework with hole DQDs in germanium heterostructures.
@@ II.1 Device
Figure 1: A double unit cell of the 2D array of spin qubit devices. The heterostructure is a 16-nm-thick Ge quantum well (red) grown on a thick Ge0.8Si0.2 buffer capped with a 50-nm-thick Ge0.8Si0.2 barrier (blue). The dot is shaped by plunger (L, R) and barrier (J, Bi) gates (gray). The yellow contour is the isodensity surface that encloses 90% of the hole charge in the ground $(1,1)$ state at bias $V_{\mathrm{L}}=V_{\mathrm{R}}=-40$ mV and $V_{\mathrm{J}}=-15$ mV (B gates grounded). The orientation of the magnetic field $\bm{B}$ is characterized by the angles $\theta$ and $\varphi$ in the crystallographic axes set $x=[100]$, $y=[010]$ and $z=[001]$.
@@ II.2 Single particle states
with $\kappa(\bm{r})=\kappa_{0}\kappa_{\mathrm{r}}(\bm{r})$, $\kappa_{0}$ the vacuum permittivity and $\kappa_{\mathrm{r}}(\bm{r})$ the dielectric constant at point $\bm{r}$. This equation is solved iteratively on a finite-volumes cartesian mesh. The bias voltages $V_{\mathrm{G}}$ on gates $\mathrm{G}\in\{\mathrm{L},\mathrm{R},\mathrm{J},\mathrm{B}_{i}\}$ are used as boundary conditions for $V_{t}(\bm{r})$. They differ from the electro-chemical potential applied on these gates by rigid shifts that depend on the ionization potentials of the materials of the gate stack and heterostructure. We next solve the Luttinger-Kohn (LK) model [109, 24, 110] for the hole wave functions $\psi_{n}$ in this potential. The physical spin components $\sigma\in\{\uparrow,\downarrow\}$ of the $\psi_{n}$’s are therefore expanded as
@@
where $u_{\nu}$ is the Bloch function with angular momentum $j_{z}=\nu$ along $z=[001]$ ($\nu=\pm 3/2$ for HH and $\nu=\pm 1/2$ for LH components), and $\varphi_{n}^{\nu}$ is an envelope function. The set of differential equations fulfilled by the $\varphi_{n}^{\nu}$’s is solved on the same mesh as the potential using finite differences (FDs). To reduce computational cost, we apply hard-wall boundary conditions ($\varphi_{n}^{\nu}=0$) $8$ nm below and $16$ nm above the Ge well. The typical mesh step in the Ge well is $\delta_{\parallel}=2$ nm along $x=[100]$ and $y=[010]$, and $\delta_{\perp}=5$ along $z$  [^1] The mesh is inhomogeneous along $z$ outside the well. We consider here holes with positive dispersion throughout this work.
Figure 2: Outline of the methodology. The $N$ lowest eigenstates of the finite-differences Luttinger-Kohn Hamiltonian are used as a basis set for fast one-particle calculations and for CI calculations. Then the lowest eigenstates of the CI Hamiltonian, sampled at different gate voltages $\{V_{\mathrm{G}}^{k}\}$ and orthogonalized against each other, are used as a “dressed” basis set for two-particle calculations over the whole operational gate voltages range.
@@
We next project $H_{1}$ in the subspace spanned by the $N$ chosen single-particle wave functions $\psi_{n}$ and introduce the $N\times N$ matrix $\hat{H}_{1}=P^{\dagger}H_{1}P$ with elements $[\hat{H}_{1}]_{ij}=\langle\psi_{i}\rvert H_{1}\lvert\psi_{j}\rangle$ (the columns of $P$ being the $\psi_{n}$’s on the FD mesh). The easiest choice for the $\psi_{n}$’s is the $N$ lowest-lying hole states at reference bias $\{V_{\mathrm{G}}^{0}\}$, in which case the matrix $\hat{H}_{0}=P^{\dagger}H_{0}P$ is diagonal. The simplest models for DQDs are typically built in such a $N=4$ basis set (the lowest two bonding and two anti-bonding states of the DQD) [55, 3]. The low-energy eigenstates of $\hat{H}_{1}$ are essentially exact at the reference bias point, and are the better the larger $N$ for finite $\delta V_{\mathrm{G}}$’s. We will discuss convergence with respect to $N$ in section II.4. The accurate description of far detuned $\delta V_{\mathrm{G}}$’s may however call for huge basis sets. To address a wider range of $\delta V_{\mathrm{G}}$’s, we may also merge the wave functions $\psi_{n}$ from $N_{\mathrm{b}}$ different $\{V_{\mathrm{G}}^{k}\}$’s in this range. The low-lying eigenstates of $\hat{H}_{1}$ are then exact at each $\{V_{\mathrm{G}}^{k}\}$ and get interpolated in between. The $\lvert\psi_{n}\rangle$ from different $\{V_{\mathrm{G}}^{k}\}$’s must be orthogonalized against each other and possibly discarded if the residual norm after orthogonalization is too small (which means that they do not bring significant information). The matrices $\hat{H}_{0}$, $\hat{D}_{\mathrm{G}}=P^{\dagger}D_{\mathrm{G}}P$, $\hat{M}_{\alpha}=P^{\dagger}M_{\alpha}P$ and $\hat{N}_{\alpha\beta}=P^{\dagger}N_{\alpha\beta}P$ are pre-computed once for all to speed up the calculation of $\hat{H}_{1}$ for arbitrary $\{\delta V_{\mathrm{G}}\}$ and $\bm{B}$.
@@ II.3 Two particle states
in the basis set of the $M=N(N-1)/2$ Slater determinants $\lvert\chi_{ij}\rangle\equiv c_{i}^{\dagger}c_{j}^{\dagger}\lvert 0\rangle$ ($i<j$). Here $c_{i}^{\dagger}$ and $c_{i}$ are the creation and annihilation operators for state $\lvert\psi_{i}\rangle$, $\lvert 0\rangle$ is the vacuum state, and the Coulomb integrals $W_{ijkl}$ read
@@
The “configuration interaction” (CI) Hamiltonian matrix $\hat{H}_{\mathrm{CI}}$ in the basis set $\lvert\chi_{ij}\rangle$ follows from straightforward commutations in Eq. (4) [80, 81]. For a single-particle basis set size $N\approx 100$, the dimension of $\hat{H}_{\mathrm{CI}}$, $M=N(N-1)/2\approx 5000$, is still inappropriate for, e.g., fast explorations of the stability diagram or time-dependent simulations. To describe the low-energy physics we are interested in, we may, therefore, want to build a reduced interacting Hamiltonian matrix $\hat{H}_{2}$ in a subspace of $M^{\prime}<M$ two-particle states $\lvert\Psi_{n}\rangle$ already dressed by Coulomb interactions. As in the single-particle case, the chosen $\Psi_{n}$’s may be the lowest-lying eigenstates of $\hat{H}_{\mathrm{CI}}$ at the reference bias $\{V_{\mathrm{G}}^{0}\}$, or a combination of such eigenstates at different $\{V_{\mathrm{G}}^{k}\}$’s (orthogonalized against each other). In that case, the low-energy eigenstates of $\hat{H}_{2}$ are “exact” at each $\{V_{\mathrm{G}}^{k}\}$ (as compared to the original $\hat{H}_{\mathrm{CI}}$). The $M^{\prime}\times M^{\prime}$ matrices of $H_{0}$, the $D_{\mathrm{G}}$’s and $M_{\alpha}$’s, and the Coulomb interaction can be computed once for all (first in the single-particle basis set $\lvert\psi_{n}\rangle$, then in the CI basis set $\lvert\chi_{ij}\rangle$ and finally in the dressed basis set $\lvert\Psi_{n}\rangle$) and combined to construct $\hat{H}_{2}$ for arbitrary magnetic fields and $\delta V_{\mathrm{G}}$’s.
As demonstrated below, we can achieve an accurate description of the DQD physics over the operational gate voltages range with as low as $N=96$ states and $M^{\prime}=112$ states. We give additional details about the implementation in Appendix A.
@@ II.4 Convergence
We illustrate the above methodology on a symmetric DQD shaped by bias voltages $V_{\mathrm{L}}=V_{\mathrm{R}}=-40$ mV with all B gates grounded.
@@
Figure 4: (a) Exchange energy $J$ as a function of $V_{\mathrm{J}}$, for different sizes of the single-particle basis set $N$ and CI basis set $M=N(N-1)/2$. The $N$ single-particle wave functions are either computed at $V_{\mathrm{J}}^{0}=V_{\mathrm{J}}$ (dots), or borrowed from a unique $V_{\mathrm{J}}^{0}=-15$ mV (lines). (b) Exchange energy $J$ at $V_{\mathrm{J}}=0$, $-5$, $-10$, $-15$, $-20$, $-25$ mV as a function of the number of single-particle states $N$ calculated at $V_{\mathrm{J}}^{0}=-15$ mV.
We next assess the convergence of the exchange energy at this symmetric operation point. At zero magnetic field, the two-particle states split into singlets and triplets. The exchange energy is defined as $J=E_{T}-E_{S}$, with $E_{S}$ and $E_{T}$ the ground-state singlet and triplet energies. $J$ is plotted as a function of $V_{\mathrm{J}}$ in Fig. 4, for different sizes of the single-particle basis set $N$ and CI basis set $M=N(N-1)/2$. The black $J(V_{\mathrm{J}})$ dots are computed with the single-particle states $\lvert\psi_{n}(V_{\mathrm{J}})\rangle$ at the same $V_{\mathrm{J}}$, while the lines are computed with a unique set of 96 $\lvert\psi_{n}(V_{\mathrm{J}}^{0})\rangle$ at $V_{\mathrm{J}}^{0}=-15$ mV. The exchange energy is found negative over a large range of gate voltages for $N=16$ (see section III), so that data are missing in the log scale plot. Although there is no “exact” reference for this quantity, we achieve good relative convergence for $N=96$ ($M=4560$) in the whole range of exchange gate voltage $-20\leq V_{\mathrm{J}}\leq-5$ mV and plunger gate detuning $|\delta V_{\mathrm{d}}|\leq 15$ mV (the latter being defined as $\delta V_{\mathrm{d}}=\delta V_{\mathrm{L}}-\delta V_{\mathrm{R}}$, with $\delta V_{\mathrm{R}}=-\delta V_{\mathrm{L}}$ in symmetric dots). In this range, the median relative variation of $E_{T}-E_{S}$ from $N=48$ to $N=96$ is $2.65\%$.
@@
Finally, we address the accuracy of the dressed two-particle basis sets. We build a basis set $\lvert\Psi_{n}\rangle$ comprising the lowest singlet and triplet eigenstates on a grid of J gate voltages $V_{\mathrm{J}}=-5$, $-10$, $-15$, and $-20$ mV and plunger gate detunings $\delta V_{\mathrm{d}}=-15$, $-10$, $-5$, 00, $5$, $10$, and $15$ mV. These $M^{\prime}=112$ $\Psi_{n}$’s are computed in the basis of $M=4560$ Slater determinants built from the $96$ single-particle wave functions at $V_{\mathrm{J}}^{0}=-15$ mV and $\delta V_{\mathrm{d}}^{0}=0$. As an illustration, the two-hole spectrum in the original $M=4560$ and dressed $M^{\prime}=112$ basis sets are plotted as a function of $\delta V_{\mathrm{d}}$ in Fig. 5 ($V_{\mathrm{J}}=-15$ meV). The mostly horizontal lines are $(1,1)$ charge states of the DQD, whereas the oblique lines are $(2,0)$ and $(0,2)$ states that anti-cross the former owing to inter-dot tunneling (see section III). The dressed basis set reproduces very well the low-lying states over the whole range of detunings. Remarkably, there are excited states missing in the dressed basis set. These are $(0,2)$ and $(2,0)$ states that do not mix with the ground $(1,1)$ states for symmetry reasons. They can not, therefore, be captured by sampling the ground singlet and triplet states in the $(\delta V_{\mathrm{d}},V_{\mathrm{J}})$ plane. However, they barely take part in exchange-driven operations in that plane  [^5] They can be slightly recoupled to $S(1,1)$ by the magnetic field (that breaks symmetries), but make negligible contributions.. The difference $\Delta_{ST}=E_{T}-E_{S}$ between the ground triplet and singlet state energies (which is the exchange $J=\Delta_{ST}$ in the $(1,1)$ sector $|\delta V_{\mathrm{d}}|\lesssim 8$ mV), as well as the relative difference between $\Delta_{ST}$ in the $M=4560$ and $M^{\prime}=112$ basis sets are plotted in the $(\delta V_{\mathrm{d}},V_{\mathrm{J}})$ plane in Fig. 6. As expected, this difference is $\approx 0$ for the points of the sampling grid, and remains typically $<1$% in-between. The dressed basis set thus matches the original CI calculation in the range $-20<V_{\mathrm{J}}<-5$ mV and $-15<\delta V_{\mathrm{d}}<15$ mV suitable for the modeling of two qubit operations, yet at a much lower numerical cost.

### v2

@@ I Introduction
In this work, we introduce a methodology for the efficient modeling of Coulomb interactions in double quantum dot (DQD) devices. Starting from reference CI calculations, we build a reduced basis set of “dressed” states accounting for the main Coulomb correlations. This basis set can describe the lowest singlet and triplet states relevant for exchange-driven operations over the whole operational gate voltages range with as few as a hundred basis functions (whereas the original CI basis set contains >4500>4500 Slater determinants). This allows for fast explorations of the device physics and for efficient time-dependent simulations. We illustrate this framework with hole DQDs in germanium heterostructures.
@@ II.1 Device
Figure 1: A double unit cell of the 2D array of spin qubit devices. The heterostructure is a 16-nm-thick Ge quantum well (red) grown on a thick Ge0.8Si0.2 buffer capped with a 50-nm-thick Ge0.8Si0.2 barrier (blue). The dots are shaped by plunger (L, R) and barrier (J, Bi) gates (gray). The yellow contour is the isodensity surface that encloses 90% of the hole charge in the ground $(1,1)$ state at bias $V_{\mathrm{L}}=V_{\mathrm{R}}=-40$ mV and $V_{\mathrm{J}}=-15$ mV (B gates grounded). The orientation of the magnetic field $\bm{B}$ is characterized by the angles $\theta$ and $\varphi$ in the crystallographic axes set $x=[100]$, $y=[010]$ and $z=[001]$.
@@ II.2 Single particle states
with $\kappa(\bm{r})=\kappa_{0}\kappa_{\mathrm{r}}(\bm{r})$, $\kappa_{0}$ the vacuum permittivity and $\kappa_{\mathrm{r}}(\bm{r})$ the dielectric constant at point $\bm{r}$. This equation is solved iteratively on a finite-volumes cartesian mesh. The bias voltages $V_{\mathrm{G}}$ on the gates $\mathrm{G}$ in the set $\{\mathrm{L},\mathrm{R},\mathrm{J},\mathrm{B}_{i}\}$ are used as boundary conditions for $V_{t}(\bm{r})$. They differ from the electro-chemical potentials applied on these gates by rigid shifts that depend on the ionization potentials of the materials of the gate stack and heterostructure. We next solve the Luttinger-Kohn (LK) model [109, 24, 110] for the hole wave functions $\psi_{n}$ in this potential (see Appendix A). The physical spin components $\sigma\in\{\uparrow,\downarrow\}$ of the $\psi_{n}$’s are therefore expanded as
@@
where $u_{\nu}$ is the Bloch function with angular momentum $j_{z}=\nu$ along $z=[001]$ ($\nu=\pm 3/2$ for HH and $\nu=\pm 1/2$ for LH components), and $\varphi_{n}^{\nu}$ is an envelope function. The set of differential equations fulfilled by the $\varphi_{n}^{\nu}$’s is solved on the same mesh as the potential using finite differences (FDs). To reduce computational cost, we apply hard-wall boundary conditions ($\varphi_{n}^{\nu}=0$) 88 nm below and $16$ nm above the Ge well. The typical mesh step in the Ge well is $\delta_{\parallel}=2$ nm along $x=[100]$ and $y=[010]$, and $\delta_{\perp}=5$ Å along $z$  [^1] The mesh is inhomogeneous along $z$ outside the well. We consider holes with positive dispersion throughout this work.
Figure 2: Outline of the methodology. The NN lowest eigenstates of the finite-differences Luttinger-Kohn Hamiltonian are used as a basis set for fast one-particle calculations and for CI calculations. Then the lowest eigenstates of the CI Hamiltonian, sampled at different gate voltages $\{V_{\mathrm{G}}^{k}\}$ and orthogonalized against each other, are used as a “dressed” basis set for two-particle calculations over the whole operational gate voltages range.
@@
We next project $H_{1}$ in the subspace spanned by the $N$ chosen single-particle wave functions $\psi_{n}$ and introduce the $N\times N$ matrix $\hat{H}_{1}=P^{\dagger}H_{1}P$ with elements $\hat{H}_{1}]_{ij}=\langle\psi_{i}\rvert H_{1}\lvert\psi_{j}\rangle$ (the columns of $P$ being the $\psi_{n}$’s on the FD mesh). The easiest choice for the $\psi_{n}$’s is the $N$ lowest-lying hole states at reference bias $\{V_{\mathrm{G}}^{0}\}$, in which case the matrix $\hat{H}_{0}=P^{\dagger}H_{0}P$ is diagonal. The simplest models for DQDs are typically built in such a $N=4$ basis set (the lowest two bonding and two anti-bonding states of the DQD) [55, 3]. The low-energy eigenstates of $\hat{H}_{1}$ are essentially exact at the reference bias point, and are the better the larger $N$ for finite $\delta V_{\mathrm{G}}$’s. We will discuss convergence with respect to $N$ in section II.4. The accurate description of far detuned $\delta V_{\mathrm{G}}$’s may however call for huge basis sets. To address a wider range of $\delta V_{\mathrm{G}}$’s, we may also merge the wave functions $\psi_{n}$ from $N_{\mathrm{b}}$ different $\{V_{\mathrm{G}}^{k}\}$’s in this range. The low-lying eigenstates of $\hat{H}_{1}$ are then exact at each $\{V_{\mathrm{G}}^{k}\}$ and get interpolated in between. The $\lvert\psi_{n}\rangle$ from different $\{V_{\mathrm{G}}^{k}\}$’s must be orthogonalized against each other and possibly discarded if the residual norm after orthogonalization is too small (which means that they do not bring significant information). The matrices $\hat{H}_{0}$, $\hat{D}_{\mathrm{G}}=P^{\dagger}D_{\mathrm{G}}P$, $\hat{M}_{\alpha}=P^{\dagger}M_{\alpha}P$ and $\hat{N}_{\alpha\beta}=P^{\dagger}N_{\alpha\beta}P$ are pre-computed once for all to speed up the calculation of $\hat{H}_{1}$ for arbitrary $\{\delta V_{\mathrm{G}}\}$ and $\bm{B}$.
@@ II.3 Two particle states
in the basis set of the M=N(N1)/2M=N(N-1)/2 Slater determinants $\lvert\chi_{ij}\rangle\equiv c_{i}^{\dagger}c_{j}^{\dagger}\lvert 0\rangle$ ($i<j$). Here $c_{i}^{\dagger}$ and $c_{i}$ are the creation and annihilation operators for state $\lvert\psi_{i}\rangle$, $\lvert 0\rangle$ is the vacuum state, and the Coulomb integrals $W_{ijkl}$ read
@@
The “configuration interaction” (CI) Hamiltonian matrix $\hat{H}_{\mathrm{CI}}$ in the basis set $\lvert\chi_{ij}\rangle$ follows from straightforward commutations in Eq. (4) [80, 81]. For a single-particle basis set size $N\approx 100$, the dimension of $\hat{H}_{\mathrm{CI}}$, $M=N(N-1)/2\approx 5000$, is still inappropriate for, e.g., fast explorations of the stability diagram or time-dependent simulations. To describe the low-energy physics we are interested in, we may, therefore, want to build a reduced interacting Hamiltonian matrix $\hat{H}_{2}$ in a subspace of $M^{\prime}<M$ two-particle states $\lvert\Psi_{n}\rangle$ already dressed by Coulomb interactions. As in the single-particle case, the chosen $\Psi_{n}$’s may be the lowest-lying eigenstates of $\hat{H}_{\mathrm{CI}}$ at the reference bias $\{V_{\mathrm{G}}^{0}\}$, or a combination of such eigenstates at different $\{V_{\mathrm{G}}^{k}\}$’s (orthogonalized against each other). In that case, the low-energy eigenstates of $\hat{H}_{2}$ are “exact” at each $\{V_{\mathrm{G}}^{k}\}$ (as compared to the original $\hat{H}_{\mathrm{CI}}$). The $M^{\prime}\times M^{\prime}$ matrices of $H_{0}$, the $D_{\mathrm{G}}$’s and $M_{\alpha}$’s, and the Coulomb interaction can be computed once for all (first in the single-particle basis set $\lvert\psi_{n}\rangle$, then in the CI basis set $\lvert\chi_{ij}\rangle$ and finally in the dressed basis set $\lvert\Psi_{n}\rangle$) and combined to construct $\hat{H}_{2}$ for arbitrary magnetic fields and $\delta V_{\mathrm{G}}$’s. A similar basis of many-body wave functions was also used for time-dependent calculations in Ref. [119].
As demonstrated below, we can achieve an accurate description of the DQD physics over the operational gate voltages range with as low as $N=96$ and $M^{\prime}=112$ states. We give additional details about the implementation in Appendix B.
@@ II.4 Convergence
We illustrate the above methodology on a symmetric DQD shaped by bias voltages $V_{\mathrm{L}}=V_{\mathrm{R}}=-40$ mV with all B gates grounded. We do not take into account the inhomogeneous strains imposed by the differential thermal contraction of materials in this section.
@@
Figure 4: (a) Exchange energy $J$ as a function of $V_{\mathrm{J}}$, for different sizes of the single-particle basis set $N$ and CI basis set M=N(N1)/2M=N(N-1)/2. The NN single-particle wave functions are either computed at $V_{\mathrm{J}}^{0}=V_{\mathrm{J}}$ (dots), or borrowed from a unique $V_{\mathrm{J}}^{0}=-15$ mV (lines). (b) Exchange energy $J$ at $V_{\mathrm{J}}=0$, 5-5, 10-10, 15-15, 20-20, 25-25 mV as a function of the number of single-particle states NN calculated at $V_{\mathrm{J}}^{0}=-15$ mV.
We next assess the convergence of the exchange energy at this symmetric operation point. At zero magnetic field, the two-particle states split into singlets and triplets. The exchange energy is defined as $J=E_{T}-E_{S}$, with $E_{S}$ and $E_{T}$ the ground-state singlet and triplet energies. $J$ is plotted as a function of $V_{\mathrm{J}}$ in Fig. 4, for different sizes of the single-particle basis set $N$ and CI basis set M=N(N1)/2M=N(N-1)/2. The black $J(V_{\mathrm{J}})$ dots are computed with the single-particle states $\lvert\psi_{n}(V_{\mathrm{J}})\rangle$ at the same $V_{\mathrm{J}}$, while the lines are computed with a unique set of 96 $\lvert\psi_{n}(V_{\mathrm{J}}^{0})\rangle$ at $V_{\mathrm{J}}^{0}=-15$ mV. The exchange energy is found negative over a large range of gate voltages for $N=16$ (see section III), so that data are missing in the log scale plot. Although there is no “exact” reference for this quantity, we achieve good relative convergence for $N=96$ ($M=4560$) in the whole range of exchange gate voltage $20\leq V_{\mathrm{J}}\leq-5$ mV and plunger gate detuning $|\delta V_{\mathrm{d}}|\leq 15$ mV (the latter being defined as $\delta V_{\mathrm{d}}=\delta V_{\mathrm{L}}-\delta V_{\mathrm{R}}$, with $\delta V_{\mathrm{R}}=-\delta V_{\mathrm{L}}$ in symmetric dots). In this range, the median relative variation of $E_{T}-E_{S}$ from $N=48$ to $N=96$ is $2.65\%$.
@@
Finally, we address the accuracy of the dressed two-particle basis sets. We build a basis set $\lvert\Psi_{n}\rangle$ comprising the lowest singlet and triplet eigenstates on a grid of J gate voltages $V_{\mathrm{J}}=-5$, 10-10, 15-15, and 20-20 mV and plunger gate detunings $\delta V_{\mathrm{d}}=-15$, 10-10, 5-5, 0, 55, $10$, and $15$ mV. These $M^{\prime}=112$ $\Psi_{n}$’s are computed in the basis of $M=4560$ Slater determinants built from the $96$ single-particle wave functions at $V_{\mathrm{J}}^{0}=-15$ mV and $\delta V_{\mathrm{d}}^{0}=0$. As an illustration, the two-hole spectrum in the original $M=4560$ and dressed $M^{\prime}=112$ basis sets are plotted as a function of $\delta V_{\mathrm{d}}$ in Fig. 5 ($V_{\mathrm{J}}=-15$ meV). The mostly horizontal lines are $(1,1)$ charge states of the DQD, whereas the oblique lines are $(2,0)$ and $(0,2)$ states that anti-cross the former owing to inter-dot tunneling (see section III). The dressed basis set reproduces very well the low-lying states over the whole range of detunings. Remarkably, there are excited states missing in the dressed basis set. These are $(0,2)$ and $(2,0)$ states that do not mix with the ground $(1,1)$ states for symmetry reasons. They can not, therefore, be captured by sampling the ground singlet and triplet states in the $(\delta V_{\mathrm{d}},V_{\mathrm{J}})$ plane. However, they hardly take part in exchange-driven operations in that plane  [^5] They can be slightly recoupled to $S(1,1)$ by the magnetic field (that breaks symmetries), but make negligible contributions.. The difference $\Delta_{ST}=E_{T}-E_{S}$ between the ground triplet and singlet state energies (which is the exchange $J=\Delta_{ST}$ in the $(1,1)$ sector $|\delta V_{\mathrm{d}}|\lesssim 8$ mV), as well as the relative difference between $\Delta_{ST}$ in the $M=4560$ and $M^{\prime}=112$ basis sets are plotted in the $(\delta V_{\mathrm{d}},V_{\mathrm{J}})$ plane in Fig. 6. As expected, this difference is $\approx 0$ for the points of the sampling grid, and remains typically <1<1% in-between. The dressed basis set thus matches the original CI calculation in the range $20<V_{\mathrm{J}}<-5$ mV and $15<\delta V_{\mathrm{d}}<15$ mV suitable for the modeling of two qubit operations, yet at a much lower numerical cost.